Find the co-ordinate of the centre of the circle and it radius if it equation is

1. x^2 +y^2 -10x-12y+6=0

2. -2X^2-2y^2-18y +9=0


1
Expert's answer
2021-06-13T15:50:51-0400

1.

"x^2+y^2-10x-12y+6=0."

"(x-5)^2+(y-6)^2-25-36+6=0."

"(x-5)^2+(y-6)^2=55."

Center: "(5,6)."

Radius: "r=\sqrt{55}."


2.

"-2x^2-2y^2-18y+9=0."

"x^2+y^2+9y-\frac{9}{2}=0."

"x^2+(y+\frac{9}{2})^2-\frac{81}{4}-\frac{9}{2}=0."

"x^2+(y+\frac{9}{2})^2=\frac{99}{4}."

Center: "(0,-\frac{9}{2})."

Radius: "r=\frac{3\sqrt{11}}{2}."